# Contemporaneous and lagged spillovers across crude oil, carbon emission allowance, climate change, and agriculture futures markets: Evidence from the $R^{2}$ decomposed connectedness approach

### v1

@@ 4.1 Averaged dynamic connectedness measures
We illustrate the averaged connectedness measures in Table 2, while values in parentheses illustrate the contemporaneous and lagged $R^{2}$ decomposed measures, respectively. We observe that on average the TCI is equal to 29.38%, indicating that on average 29.38% of variance of the LHS variables can be explained by the RHS variables. Decomposing this metric into its contemporaneous and lagged components, we find that 20.43% are attributed to contemporaneous effects while only 8.95% are caused by lagged dynamics. Furthermore, almost all the contemporaneous FROM and TO connectedness metrics are significantly greater than the corresponding lagged counterparts. However, cocoa, milk, and barley have lagged FROM measures (11.03%, 9.81%, and 13.34%) exceeding their FROM contemporaneous measures (10.28%, 7.65%, and 10.84%). The lagged TO measures for cattle and milk (10.14% and 7.29%, respectively) are both higher than their contemporaneous values (9.45% and 6.89%, respectively).
Additionally, we observe that corn is the primary net transmitter of shocks (7.01%), followed by soybean meal (5.59%), and wheat (3.51%). Several factors might make corn the main net contributor of shocks. First, the close relationship between corn and commodity markets has been widely reported (Kang et al., 2017). Corn is one of the primary raw materials for biofuels, whose demand is closely linked to oil prices. Secondly, corn, requiring a significant amount of water, is particularly sensitive to climate change. Third, corn plays a crucial role in the global food supply chain. Moreover, agricultural activities can also lead to significant carbon emission allowance. Similarly, both soybean meal and wheat are important agricultural products for many countries, which might make them risk contributors. Conversely, barley is the chief net receiver of shocks (-7.73%), with cocoa (-3.68%) and leah hog (-3.59%) following. Besides that, we observe that climate change is a net transmitter (1.57%), while crude oil (-0.22%) and carbon emission allowance (-1.30%) are net risk receivers. In detail, carbon emission (2.12%), cattle (1.99%), and milk (2.16%) receive the largest risk from climate change, while the largest spillovers to climate change are caused by crude oil (3.1%). Barley receives the largest spillovers from wheat (3.95%), with 0.84% being caused by contemporaneous effects while 3.11% originate from lagged spillovers. More specifically, corn and wheat receive major spillovers from each other. The spillover among crude oil and agricultural markets is consistent with the findings of Du et al. (2011).
@@ 4.2 Dynamic total connectedness
Fig. 1 presents the dynamic total connectedness. We note high market spillovers since 2021. Subsequently, the dynamic total connectedness continued to decrease, implying that the impact of the COVID-19 pandemic is gradually diminishing. However, early in the outbreak of the Russia-Ukraine war, the dynamic total connectedness increased significantly. Afterward, a subsequent decline is observed, with a peak of 28% in June 2023. Since then, the TCI value has reached a plateau of around 25% throughout the period of investigation. Overall, the lagged TCI remained relatively constant except for its increase during the outbreak of the Russia-Ukraine war, which fluctuated around 10% during the entire sample period.
Moreover, we see that the overall and contemporaneous dynamic total connectedness measures behaved extremely similarly. Hence, the contemporaneous interdependencies are more significant than the lagged interdependencies, as they account for the majority of total risk dynamics. The movement in the overall TCI is mainly driven by the contemporaneous dynamics and not by the lagged dynamics, which is in line with the work of Balli et al. (2023).
@@ 4.3 Dynamic FROM, TO and NET directional connectedness
Figures 2 and 3 illustrate the FROM and TO directional connectedness, respectively. We find that the contemporaneous spillovers are more pronounced than the lagged spillovers. In Fig. 2, the FROM connectedness for most variables experienced significant fluctuations during the Russia-Ukraine war, especially the crude oil, climate change, wheat and barley. Additionally, soybean has the highest degree of FROM connectedness, followed by soybean oil and wheat.
@@
In addition, a net positive spillover corresponds to a net transmitter of risks to other markets, while a net negative spillover suggests that a particular market is a net receiver of shocks in the system. Fig. 4 presents that the net contributors of shocks in the systems during the sample period mainly include climate change, corn, wheat, oats, soybean meat, and soybean oil, while cocoa, lean hog, juice, and barley mainly act as risk receivers of shocks from other variables. Barley is the biggest receiver of net spillovers. The remaining markets at different periods act as transmitters and receivers alternatively.
A critical look suggests that crude oil behaved as a receiver then switched roles as a transmitter since November 2022. However, carbon emission allowance shows an almost opposite pattern. Moreover, lagged net spillover of most markets exhibits higher similarity to overall net spillover effects, which is in line with Liu et al. (2024). However, different from other markets, soybean initially displayed a higher level of contemporaneous spillovers, followed by a stronger level of lagged correlation.
@@
4.4 Net pairwise directional connectedness measures
In this subsection, we discuss how one specific market impacts another on a pairwise basis. Fig. 5 shows the net pairwise risk spillovers of the system. By taking a closer look, we find obvious spillovers between climate change and other variables. In the early stages of our selected sample period, climate risk has a net transmitting effect on the crude oil market, which is consistent with Jin et al. (2023). Cocoa, oats, and milk are the recipients of spillover from climate change. On the other hand, climate change receives spillover from the following sectors: corn and soybean oil.
Moreover, we notice that crude oil is a recipient of shocks from soybean, soybean meat, and a contributor of shocks to rice, cotton, and coffee. However, this result differs from the findings of Tiwari et al. (2022). And the net pairwise directional connectedness between crude oil and sugar switches its sign during the sample period, which corroborates the work of Tiwari et al. (2022).
Additionally, we find that lean hog is constantly dominated by soybean, corn, oats, soybean meal, and barley, which accords with Table 2. And barley receives high levels of spillovers from carbon emission allowance, wheat, and corn. Meanwhile, corn and rice transfer risks to the carbon market. We also observe spillovers from cattle, soybean, and soybean meal to rice.
@@ 4.5 Net pairwise network analysis
In Fig. 6, we visually illustrate the net pairwise directional connectedness measures. A transmitter is denoted by the blue dot, while a receiver is denoted by the yellow dot. Carbon emission allowance, barley, and rice exhibit negative net spillover effects. Corn and wheat have notable positive risk spillover effects, which is in line with Guo et al. (2023). Moreover, there is a high level of overall and lagged connectedness in wheat and barley. This strong correlation aligns with Banerjee et al. (2024).
@@ 4.6 Robustness check
Fig. 7 illustrates the robustness checks based on the dynamic total connectedness. We replace the Pearson correlation coefficients with the Spearman and Kendall correlation coefficients. The dynamic TCIs of the three kinds of correlation coefficients are very close, implying that the empirical results are robust.
Figure 7: Robustness check: Dynamic total connectedness. The black line depicts the overall dynamic total connectedness while the dynamic contemporaneous and lagged connectedness are illustrated in red and green, respectively. The three distinct line types represent three different correlation coefficients: Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$.
@@ 5 Conclusion
Climate change has attracted global attention, with profound impacts on agricultural activities, energy extraction, and transportation. To promote a low-carbon economy, protect the environment, and maintain financial stability, it is essential to understand the risk spillovers between climate change, crude oil, carbon emission allowance, and agricultural markets. In this paper, we utilize a novel $R^{2}$ decomposed connectedness approach to investigate the contemporaneous and lagged connectedness across these markets.
The analysis reveals that the total TCI dynamics have been mainly driven by contemporaneous dynamics, which are more volatile than the lagged dynamics. Moreover, the total connectedness has increased due to the Russia-Ukrainian war. Besides that, there is a degree of spillover from climate change to other markets. In particular, the largest transmitter of carbon emission allowance, cattle, and milk is climate change. And there exist heterogenous spillover effects among agricultural markets. In detail, corn is the primary net transmitter of shocks, while barley mainly acts as a net receiver.
Our results have the following implications. By distinguishing between contemporaneous and lagged spillover effects, investors and risk managers can more accurately assess market risks, identify sources of risk and transmission pathways, and develop more effective risk management strategies. The detailed analysis enables market participants to predict future market changes in advance, adjust their investment strategies in a timely manner, and enhance market responsiveness.
@@ Acknowledgements
This work was supported by the Shanghai Planning Office of Philosophy and Social Science (2022EJB006).

### v2

@@ 4.1 Averaged dynamic connectedness measures
Table 2 shows the averaged connectedness measures, with parentheses indicating contemporaneous and lagged $R^{2}$ decomposed values. The average TCI is 29.38%, with 20.43% from contemporaneous effects and 8.95% from lagged dynamics, indicating that 29.38% of the LHS variables’ variance is explained by the RHS variables.
Furthermore, almost all the contemporaneous FROM and TO connectedness metrics are significantly greater than the corresponding lagged counterparts. However, cocoa, milk, and barley have lagged FROM measures (11.03%, 9.81%, and 13.34%) exceeding their FROM contemporaneous measures (10.28%, 7.65%, and 10.84%). The lagged TO measures for cattle and milk (10.14% and 7.29%, respectively) are both higher than their contemporaneous values (9.45% and 6.89%, respectively).
@@
Additionally, corn is the primary net transmitter of shocks (7.01%), followed by soybean meal (5.59%), and wheat (3.51%). Several factors might make corn the main net contributor of shocks. First, the close relationship between corn and commodity markets has been widely reported (Kang et al., 2017). Corn is one of the primary raw materials for biofuels, whose demand is closely linked to oil prices. Secondly, corn, requiring a significant amount of water, is particularly sensitive to climate change. Third, corn plays a crucial role in the global food supply chain. Moreover, agricultural activities can also lead to significant carbon emission allowance. Similarly, both soybean meal and wheat are important agricultural products for many countries, which might make them risk contributors.
Conversely, barley is the chief net receiver of shocks (-7.73%), with cocoa (-3.68%) and leah hog (-3.59%) following. Besides that, climate change is a net transmitter (1.57%), while crude oil (-0.22%) and carbon emission allowance (-1.30%) are net risk receivers. In detail, carbon emission (2.12%), cattle (1.99%), and milk (2.16%) receive the largest risk from climate change, while the largest spillovers to climate change are caused by crude oil (3.1%). Barley receives the largest spillovers from wheat (3.95%), with 0.84% being caused by contemporaneous effects while 3.11% originate from lagged spillovers. More specifically, corn and wheat receive major spillovers from each other. The spillover between crude oil and agricultural markets aligns with Du et al. (2011).
@@ 4.2 Dynamic total connectedness
Fig. 3 presents the dynamic total connectedness. We note high market spillovers since 2021. The dynamic total connectedness decreased, indicating a diminishing COVID-19 impact, but rose sharply early in the Russia-Ukraine war. A decline followed, peaking at 28% in June 2023. Then the TCI has plateaued at around 25% throughout the investigation period. The lagged TCI stayed around 10%, with an increase during the Russia-Ukraine war.
@@
Moreover, the overall and contemporaneous dynamic total connectedness measures behaved extremely similarly. Contemporaneous interdependencies dominate lagged ones, driving most total risk dynamics. The movement in the overall TCI is mainly driven by the contemporaneous dynamics and not by the lagged dynamics, which is in line with the work of Balli et al. (2023).
@@ 4.3 Dynamic FROM, TO and NET directional connectedness
Fig. 4 and Fig. 5 illustrate the FROM and TO directional connectedness, respectively. The contemporaneous spillovers are more pronounced than the lagged spillovers. In Fig. 4, most variables’ FROM connectedness fluctuated significantly during the Russia-Ukraine war, especially the crude oil, climate change, wheat and barley. Additionally, soybean has the highest degree of FROM connectedness, followed by soybean oil and wheat.
@@
Fig. 6 presents that the net contributors of shocks mainly include climate change, corn, wheat, oats, soybean meat, and soybean oil, while cocoa, lean hog, juice, and barley mainly act as risk receivers. Among them, barley is the biggest receiver of net spillovers. The remaining markets alternate between being transmitters and receivers.
A critical look suggests that crude oil shifts from a receiver to a transmitter since November 2022. However, carbon emission allowance shows an almost opposite pattern. Moreover, lagged net spillover of most markets exhibits higher similarity to overall net spillover effects, in line with Liu et al. (2024). Unlike other markets, soybean shows higher contemporaneous spillovers, followed by stronger lagged correlations.
4.4 Net pairwise directional connectedness measures
Fig. 7 shows the net pairwise risk spillovers of the system. A closer look reveals clear spillovers between climate change and other variables. Early in the sample period, climate risk is a net transmitter to the crude oil market, consistent with Jin et al. (2023). Cocoa, oats, and milk receive spillovers from climate change. Meanwhile, climate change receives spillovers from corn and soybean oil.
Additionally, crude oil receives shocks from soybean and soybean meal, while contributing to rice, cotton, and coffee. However, this result differs from Tiwari et al. (2022). And the net pairwise directional connectedness between crude oil and sugar switches its sign during the sample period, which corroborates the work of Tiwari et al. (2022).
@@ 4.4 Net pairwise directional connectedness measures
Moreover, lean hog is constantly dominated by soybean, corn, oats, soybean meal, and barley, which accords with Table 2. And barley receives high levels of spillovers from carbon emission allowance, wheat, and corn. Meanwhile, corn and rice transfer risks to the carbon market. We also observe spillovers from cattle, soybean, and soybean meal to rice.
@@ 4.5 Net pairwise network analysis
Fig. 8 illustrates the net pairwise directional connectedness measures, with blue dots for transmitters and yellow dots for receivers. Carbon emission allowance, barley, and rice exhibit negative net spillover effects. Corn and wheat have notable positive risk spillover effects, which is in line with Guo et al. (2023). Moreover, there is a high level of overall and lagged connectedness in wheat and barley, aligning with Banerjee et al. (2024).
@@ 4.6 Robustness check
Fig. 9 illustrates the robustness checks based on the dynamic total connectedness. We replace Pearson with Spearman and Kendall correlation coefficients and use the DY method as an alternative to verify TCI robustness. The dynamic TCIs of the three correlation types are similar, indicating robust results.
Figure 9: Robustness check: Dynamic total connectedness. The black line depicts the overall dynamic total connectedness while the dynamic contemporaneous and lagged connectedness are illustrated in red and green, respectively. The three distinct line types represent three different correlation coefficients: Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$. The blue line illustrates the DY method.
@@ 5 Conclusion
This paper utilizes a novel $R^{2}$ decomposed connectedness approach to identify the contemporaneous and lagged effects across futures markets of agriculture, crude oil, carbon emission allowance, and climate change. The analysis shows the total TCI is driven by more volatile contemporaneous dynamics than lagged ones. Moreover, the total connectedness has increased due to the Russia-Ukrainian war. Additionally, there exist heterogenous spillover effects among agricultural markets. In detail, corn, soybean meal and wheat are the primary net transmitters, while barley, cocoa and lean hog act as main risk receivers. And we observe a degree of spillover from climate change to other markets. As our study focuses on the current factors influencing agricultural market spillovers, future research could explore the impact of fintech and other financial innovations on agricultural market risk transmission.
Our results have the following implications: Given that contemporaneous spillover effects are more significant than lagged ones, policymakers and investors should focus on real-time monitoring and swift response mechanisms to manage risks as they occur. Moreover, they should implement differentiated risk management measures for various agricultural products, especially key commodities. In detail, hedging strategies could be more aggressive for contributors, while protective measures might be prioritized for receivers. Additionally, policymakers should invest in risk-resilient agricultural practices, strengthening responses to multiple risk sources.
@@ Acknowledgements
The authors would like to acknowledge the support of the Shanghai Planning Office of Philosophy and Social Science (2022EJB006).